# Accessible, At-Home Detection of Parkinson’s Disease via Multi-task Video Analysis

### v3

@@ Results
AUROC AUPRC $\mathbf{F_{1}}$ score
PPV
NPV
Sensitivity
Specificity Coverage
Majority Voting 85.3 83.9 89.6 78.0 78.2 80.0 87.9 76.5 89.9 -
Neural Late Fusion $84.1\pm 0.4$ $81.3\pm 4.8$ $91.7\pm 2.2$ $\mathbf{86.7\pm 3.1}$ $73.2\pm 8.3$ $73.5\pm 7.5$ $89.2\pm 3.3$ $76.3\pm 9.4$ $88.2\pm 2.7$
Early Fusion Baseline $83.6\pm 0.6$ $81.8\pm 0.7$ $91.0\pm 0.2$ $85.8\pm 0.4$ $76.7\pm 0.7$ $75.4\pm 1.1$ $88.3\pm 0.4$ $78.1\pm 0.9$ $86.5\pm 0.8$
Hybrid Fusion Baseline $84.1\pm 0.3$ $82.4\pm 0.4$ $91.4\pm 0.2$ $86.5\pm 0.3$ $77.3\pm 0.4$ $76.2\pm 0.7$ $88.5\pm 0.3$ $78.6\pm 0.6$ $87.0\pm 0.6$
UFNet - Early Fusion $86.7\pm 0.5$ $85.8\pm 0.5$ $92.7\pm 0.3$ $86.2\pm 0.5$ $79.9\pm 0.8$ $83.3\pm 0.7$ $88.3\pm 0.6$ $76.9\pm 1.4$ $91.9\pm 0.4$
UFNet - Early Fusion
(withhold uncertain preds)
$87.5\pm 0.4$ $86.5\pm 0.5$ $92.9\pm 0.3$ $86.4\pm 0.7$ $80.7\pm 0.8$ $84.0\pm 0.7$ $89.1\pm 0.6$ $77.7\pm 1.4$ $92.4\pm 0.4$ $97.4\pm 0.3$
UFNet - Hybrid Fusion $87.3\pm 0.4$ $86.4\pm 0.4$ $92.8\pm 0.2$ $86.3\pm 0.5$ $81.0\pm 0.6$ $83.8\pm 0.5$ $89.0\pm 0.4$ $78.4\pm 1.0$ $92.0\pm 0.3$
UFNet - Hybrid Fusion
(withhold uncertain preds)
$\mathbf{88.0\pm 0.3}$ $\mathbf{87.1\pm 0.3}$ $\mathbf{93.0\pm 0.2}$ $86.5\pm 0.5$ $\mathbf{81.8\pm 0.6}$ $\mathbf{84.6\pm 0.5}$ $\mathbf{89.7\pm 0.4}$ $\mathbf{79.3\pm 0.9}$ $\mathbf{92.6\pm 0.2}$ $97.8\pm 0.3$
Table 4: Comparison of UFNet performance against multimodal baselines. The underlined metrics indicate significantly better performance compared to all four baselines. Bold metrics indicate overall best performance across all choices. All scores should be interpreted as percentages (%). Confidence interval is not reported for majority voting since it does not involve any randomness.
(a) ROC curve.
(b) Model calibration.
Figure 2: ROC and calibration curves showing the performance of the best models (with different random seeds). The shaded regions and the error bars represent 95% confidence intervals.
Comparison against baselines
The proposed uncertainty-calibrated fusion network (UFNet) outperformed all four baseline methods in most metrics (see Table 4). While the neural late fusion baseline achieved the best AUPRC, its average $F_{1}$ score is significantly lower compared to the other baselines and UFNet models. This baseline also demonstrates the least stability, with a large confidence interval for most metrics. UFNet significantly improves accuracy, balanced accuracy, PPV, and specificity over the baselines, with other metrics being non-inferior.
Ablation results
Including task-specific predictions as additional input (hybrid fusion) slightly enhanced model performance. Withholding uncertain predictions also boosted performance. With dropped uncertain predictions, the best UFNet (hybrid fusion) model achieved $88.0\pm 0.3\%$ accuracy, $87.1\pm 0.3\%$ balanced accuracy, $93.0\pm 0.2\%$ AUROC, and $81.8\pm 0.5\%$ $F_{1}$ score, outperforming all other models. It also achieves the best PPV ($84.6\pm 0.5\%$), NPV ($89.7\pm 0.4\%$), sensitivity ($79.3\pm 0.9\%$), and specificity ($92.6\pm 0.3\%$). However, the model can now predict for $97.8\pm 0.3\%$ of the data where it is certain enough. In addition, the expected calibration error (ECE) of the model is $5.4\pm 0.5\%$, and the Brier score is $0.097\pm 0.002$, indicating that the model predicted probability is aligned with true disease probability (Figure 2).
Figure 3: Misclassification rate of the best UFNet model across demographic subgroups. The analysis is done on the test set participants. Error bars demonstrate 95% confidence intervals.
Performance across demographic subgroups
No significant bias in model performance was observed based on sex or race in the test set (162 individuals). The average error (i.e., misclassification) rate across female participants ($n=85$) was $14.1\pm 7.4\%$, compared to $6.5\pm 5.5\%$ for male participants ($n=77$). This difference, though notable, is not statistically significant (p-value = 0.11). The error rate was $7.63\pm 4.79\%$ for white participants ($n=118$) and $5.56\pm 11.39\%$ for non-white participants ($n=18$). Based on Fisher’s exact test, this difference was also non-significant (Fisher’s odd ratio = 0.71, p-value = 1.0). However, the error rate varied notably based on age subgroups. The model performed well for individuals aged 50 to 80, while the error rate was higher for those aged 30 to 50 or over 80, likely due to under-representation of these age groups, as the dataset primarily consists of 50-80 year old individuals (77.1% of the entire dataset).
Discussions and Limitations
Video analysis provides an accessible, cost-effective, and convenient means of screening for PD, particularly benefiting individuals in remote areas or low-income countries with limited access to neurological care. Building on prior work demonstrating the feasibility of classifying PD symptoms from single video tasks (Rahman et al. 2021; Islam et al. 2023; Adnan et al. 2023), this study explores a combination of three tasks for a more holistic and generalizable approach, reflecting the multifaceted nature of PD symptoms. We developed models to classify PD cases with high accuracy, sensitivity, and specificity using webcam recordings of individuals performing finger tapping, speech, and smile tasks.
We carefully selected three proposed tasks for remote completion, considering feasibility and safety. These tasks assess bradykinesia, hypomimia, and speech impairment and can be safely performed at home without assistance. While gait analysis is common for evaluating PD (Li and Li 2022; Liu et al. 2022a), it presents logistical challenges and potential risks of falling for PD patients. In contrast, the finger-tapping task offers a safer alternative for motor assessment. Facial expressions and speech are assessed through natural conversations, but prompting natural conversations introduces subjectivity and confounding factors, hindering machine learning model training with limited data. Mimicking a smile provides a standardized alternative, and the pangram utterance task reduces confounding factors in speech analysis. While sustained phonation (holding a vowel sound for as long as possible) is another speech assessment option (Tsanas et al. 2009; Vaiciukynas et al. 2017), its analysis becomes complicated due to inconsistencies across various recording devices. All selected tasks demonstrated reasonable predictive performance in differentiating individuals with and without PD, validating our task selection.
Models trained on all three tasks performed better than single-task models. Among single-task models, PD classification from the speech task performed the best. However, the UFNet model trained on all three tasks was our best-performing model, achieving the highest AUROC, accuracy, recall, and specificity. Withholding uncertain predictions further boosted the model’s performance, ensuring additional safeguards against potential miss-predictions. The proposed model also outperformed traditional data-efficient neural models for combining multiple modalities.
However, our work has limitations. The model performs consistently across sex and race subgroups, but accuracy drops for younger ($30-49$) and older (above 80) age groups. We find that participants from these age groups are under-represented in our dataset (see Technical Appendix A containing demographic summary), with over 75% of the participants falling within the 50-80 age range. This age bias likely explains the model’s more robust performance in this middle range. It’s also important to note that younger individuals are less likely to have PD, as it is more commonly associated with aging. Until a more balanced dataset is available, we recommend applying the tool only for individuals between 50 and 80 years old. Future work should prioritize recruiting younger and older participants.
Additionally, the thresholds used in the study could be customized based on individual preferences, as the decision threshold directly affects the model’s sensitivity and specificity. We used a common 0.5 threshold, but individual preferences for risk-benefit trade-offs might necessitate adjustments. For instance, some users might seek clinical evaluation even at lower probabilities (i.e., prefer high sensitivity), while others might wait for a higher likelihood before incurring healthcare costs (i.e., prefer high specificity). We chose to withhold predictions when the model is uncertain about the PD/non-PD classification. If the 95% confidence interval of the prediction contains both positive and negative classes, the model refrains from making a prediction. In the absence of MC-dropout, an alternative is interpreting the model prediction scores as its confidence (e.g., a model is more confident when the prediction score is $0.95$ than a score of $0.55$) and setting a suitable confidence level based on accuracy-coverage trade-off and participant preference. Other alternatives for withholding predictions can also be explored in the future.
Furthermore, as videos are primarily gathered in an unsupervised fashion, issues such as noncompliance with task instructions and various forms of noise are common. For example, during the finger-tapping task, many participants performed fewer than the required ten taps, often with their task-performing hand obscured from view. Background noise may distort speech features, and the presence of multiple individuals in the frame could compromise smile feature extraction. Integrating post-hoc quality assessment algorithms into the data collection framework could enhance data quality by identifying quality issues and prompting the user to re-record tasks if needed.
A comprehensive, large-scale dataset for Parkinson’s disease is currently unavailable due to the lack of substantial resources and clinical collaboration. This paper addresses this gap by introducing the largest multi-task video dataset in the literature and publicly releasing the anonymized features. Unfortunately, we could not gather multiple datasets of videos or extract similar features to compare model performance across different datasets since patient videos, being protected health information, are not publicly accessible. Making our dataset available can help bridge this gap, facilitating broader comparisons and furthering research in this area.
Conclusion

### v4

@@ Appendix A. Supplementary Results
$F_{1}$ Score AUROC AUPRC
Custom features $\mathbf{77.6\pm 0.2}$ $\mathbf{74.4\pm 0.2}$ $\mathbf{67.5\pm 0.3}$ $\mathbf{83.6\pm 0.1}$ $\mathbf{65.4\pm 0.1}$
VideoMAE $73.5\pm 0.4$ $69.1\pm 0.8$ $43.2\pm 2.1$ $67.3\pm 0.5$ $52.5\pm 0.7$
Uniformer $75.7\pm 0.4$ $71.1\pm 0.6$ $53.8\pm 0.7$ $74.9\pm 0.4$ $59.9\pm 0.3$
ViViT $74.5\pm 0.5$ $72.2\pm 1.0$ $41.6\pm 3.0$ $73.9\pm 0.4$ $58.3\pm 0.8$
TimeSformer $73.9\pm 0.4$ $69.4\pm 0.6$ $49.1\pm 1.9$ $75.5\pm 0.3$ $58.9\pm 0.5$
Table 11: Smile task performance
Choosing decision threshold
Figure 3: Effect of decision threshold. For this plot, we used the best performing UFNet model with 30 different random seeds. Shaded regions show the 95% confidence intervals.
Although it is common practice to choose 0.50 as the threshold for binary classification, this can be further customized based on individual preference and underlying healthcare infrastructure. Here we explore how different decision thresholds impacts the sensitivity and specificity of the proposed UFNet model (see Figure 3) on the test set. As expected, with a higher decision threshold, the specificity of the model notably improves, but at the expense of reduced sensitivity. Likewise, a lower decision threshold yields higher sensitivity while the specificity gets penalized. In resource-constrained environments, a higher threshold might be preferred to minimize false negatives. Conversely, when healthcare resources are abundant, a lower threshold can prioritize identifying all potential cases.
Performance Analysis across Disease Duration
To evaluate the model’s performance across various stages of PD, we analyzed its misclassification rates relative to disease duration. From 102 participants with duration data, 15 were included in the test set (25 independent samples) covering 10 unique duration values. An initial exploratory analysis using Kendall’s Tau revealed a weak negative correlation between disease duration and misclassification rates (Tau = -0.33, p = 0.39), suggesting a slight trend of decreasing error rates with longer disease duration. However, these findings were not statistically significant due to the small sample size.
Appendix B. Hyper-parameter search for the task-specific models
Task-specific models without Monte Carlo dropout
The hyper-parameter search space is outlined in Table 12. The selected hyper-parameters for the task-specific models are mentioned below:
Hyperparameter Values/range Distribution
batch size {256, 512, 1024} Categorical
learning rate [0.0005, 1.0] Uniform
drop correlated features? {“yes”, “no”} Categorical
correlation threshold {0.80, 0.85, 0.90, 0.95} Categorical
use feature scaling? {“yes”, “no”} Categorical
scaling method {“StandardScaler”, “MinMaxScaler”} Categorical
use minority oversampling (i.e., SMOTE)? {“yes”, “no”} Categorical
number of hidden layers {0, 1} Categorical
number of epochs [1, 100] Uniform Integer
optimizer {“SGD”, “AdamW”} Categorical
momentum [0.1, 1.0] Uniform
use scheduler? {“yes”, “no”} Categorical
scheduler {“step”, “reduce on plateau”} Categorical
step size [1, 30] Uniform Integer
patience [1, 20] Uniform Integer
gamma [0.5, 0.95] Uniform
seed [100, 999] Uniform
Table 12: Hyper-parameter search space for the task-specific models (without MC dropout).
Finger-tapping task (both hands):
batch size = 256, learning rate = 0.6246956232061768, drop correlated features? = no, use feature scaling? = yes, scaling method = StandardScaler, use minority oversampling? = no, number of hidden layers = 0, number of epochs = 82, optimizer = SGD, momentum = 0.8046223742478498, use scheduler? = no, seed = 276
Finger-tapping task (left hand):
batch size = 512, learning rate = 0.807750048295928, drop correlated features? = yes, correlation threshold = 0.95, use feature scaling? = yes, scaling method = StandardScaler, use minority oversampling? = no, number of hidden layers = 0, number of epochs = 50, optimizer = SGD, momentum = 0.6614402107331798, use scheduler? = no, seed = 556
Finger-tapping task (right hand):
batch size = 512, learning rate = 0.5437653223933676, drop correlated features? = no, use feature scaling? = yes, scaling method = StandardScaler, use minority oversampling? = no, number of hidden layers = 1, number of epochs = 74, optimizer = SGD, momentum = 0.709095892070382, use scheduler? = no, seed = 751
Smile:
batch size = 1024, learning rate = 0.8365099039036598, drop correlated features? = no, use feature scaling? = yes, scaling method = StandardScaler, use minority oversampling? = yes, number of hidden layers = 0, number of epochs = 74, optimizer = SGD, momentum = 0.615229008837764, use scheduler? = yes, scheduler = reduce on plateau, patience = 4, seed = 488
Speech:
batch size = 256, learning rate = 0.06573643554880117, drop correlated features? = no, use feature scaling? = yes, scaling method = StandardScaler, use minority oversampling? = no, number of hidden layers = 1, number of epochs = 27, optimizer = SGD, momentum = 0.5231696483982686, use scheduler? = no, seed = 287
Task-specific models with Monte Carlo dropout
The hyper-parameter search space is outlined in Table 13. The selected hyper-parameters for the task-specific models are mentioned below:
Hyperparameter Values/range Distribution
batch size {256, 512, 1024} Categorical
learning rate [0.0005, 1.0] Uniform
drop correlated features? {“yes”, “no”} Categorical
correlation threshold {0.80, 0.85, 0.90, 0.95} Categorical
use feature scaling? {“yes”, “no”} Categorical
scaling method {“StandardScaler”, “MinMaxScaler”} Categorical
use minority oversampling (i.e., SMOTE)? {“yes”, “no”} Categorical
number of hidden layers {0, 1} Categorical
dropout probability [0.01, 0.30] Uniform
number MC dropout rounds {100, 300, 500, 1000, 3000, 5000, 10000} Categorical
number of epochs [1, 100] Uniform Integer
optimizer {“SGD”, “AdamW”} Categorical
momentum [0.1, 1.0] Uniform
use scheduler? {“yes”, “no”} Categorical
scheduler {“step”, “reduce on plateau”} Categorical
step size [1, 30] Uniform Integer
patience [1, 20] Uniform Integer
gamma [0.5, 0.95] Uniform
Table 13: Hyper-parameter search space for the task-specific models (with MC dropout).
Finger-tapping task:
batch size = 256, learning rate = 0.3081959128766984, drop correlated features? = no, use feature scaling? = yes, scaling method = StandardScaler, use minority oversampling? = no, number of hidden layers = 0, dropout probability = 0.24180259124462203, number of MC dropout rounds = 1000, number of epochs = 87, optimizer = SGD, momentum = 0.9206317439937552, use scheduler? = yes, scheduler = reduce on plateau, patience = 13, seed = 790
Smile task:
batch size = 256, learning rate = 0.03265227174722892, drop correlated features? = no, use feature scaling? = yes, scaling method = StandardScaler, use minority oversampling? = yes, number of hidden layers = 0, dropout probability = 0.10661756438565197, number of MC dropout rounds = 1000, number of epochs = 64, optimizer = SGD, momentum = 0.5450637936769563, use scheduler? = no, seed = 462
Speech task:
batch size = 1024, learning rate = 0.364654919080181, drop correlated features? = yes, correlation threshold = 0.95, use feature scaling? = no, use minority oversampling? = no, number of hidden layers = 0, dropout probability = 0.23420212038821583, number of MC dropout rounds = 10000, number of epochs = 74, optimizer = AdamW, use scheduler? = no, seed = 303
Appendix C. Hyper-parameter search for the UFNet models
The hyper-parameter search space is outlined in Table 14. The selected hyper-parameters are mentioned below:
Hyperparameter Values/range Distribution
batch size {$256,512,1024$} Categorical
learning rate [$1.05e^{-5},$ 1.0] Uniform
use minority oversampling? {“yes”, “no”} Categorical
oversampling method
{“SMOTE”, “SVMSMOTE”, “ADASYN”,
“BoarderlineSMOTE”, “SMOTEN”,
“RandomOversampler”}
Categorical
number of hidden layers {$1$} Categorical
projection dimension {$128,256,512$} Categorical
query (query/key/value) dimension {$32,64,128,256$} Categorical
hidden dimension {$4,8,16,32,64,128$} Categorical
dropout probability [$0.05,0.50$] Uniform
$\eta$ [$0.1,100$] Uniform
number MC dropout rounds {$30$} Categorical
number of epochs [$1,300$] Uniform Integer
optimizer {“SGD”, “AdamW”, “RMSprop”} Categorical
momentum [$0.1,1.0$] Uniform
use scheduler? {“yes”, “no”} Categorical
scheduler {“step”, “reduce on plateau”} Categorical
step size [$1,30$] Uniform Integer
patience [$1,20$] Uniform Integer
gamma [$0.5,0.95$] Uniform
Table 14: Hyper-parameter search space for the UFNet models.
Finger-tapping + Smile + Speech: batch size = 1024, learning rate = 0.020724, use minority oversampling? = no, number of hidden layers = 1, projection dimension = 512, query dimension = 64, hidden dimension = 128, dropout probability = 0.4959892, $\eta$ = 81.8179035, number of MC dropout rounds = 30, number of epochs = 164, optimizer = SGD, momentum = 0.689782158, use scheduler? = no, seed=242
Finger-tapping + Smile: batch size = 256, learning rate = 0.06754950185131235, use minority oversampling? = no, number of hidden layers = 1, projection dimension = 512, query dimension = 64, hidden dimension = 64, dropout probability = 0.4453733432524283, $\eta$ = 12.554916213821272, number of MC dropout rounds = 30, number of epochs = 18, optimizer = SGD, momentum = 0.9822830376765904, use scheduler? = yes, scheduler = reduce on plateau, patience = 10, seed=919
Finger-tapping + Speech: batch size = 512, learning rate = 0.04035092571261426, use minority oversampling? = no, number of hidden layers = 1, projection dimension = 256, query dimension = 256, hidden dimension = 16, dropout probability = 0.49813214914563847, $\eta$ = 79.95872101951133, number of MC dropout rounds = 30, number of epochs = 164, optimizer = SGD, momentum = 0.24020164138826405, use scheduler? = yes, scheduler = reduce on plateau, patience = 12, seed=953
Smile + Speech: batch size = 512, learning rate = 0.16688970966723005, use minority oversampling? = no, number of hidden layers = 1, projection dimension = 128, query dimension = 64, hidden dimension = 4, dropout probability = 0.3763157755397192, $\eta$ = 51.88439832518041, number of MC dropout rounds = 30, number of epochs = 132, optimizer = SGD, momentum = 0.22419387711544064, use scheduler? = yes, scheduler = reduce on plateau, patience = 13, seed=845
Reproducibility Checklist
  1. 1.
    This paper:
    • Includes a conceptual outline and/or pseudocode description of AI methods introduced – yes
    • Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results – yes
    • Provides well marked pedagogical references for less-familiare readers to gain background necessary to replicate the paper – yes
  2. 2.
    Does this paper make theoretical contributions? – no
  3. 3.
    Does this paper rely on one or more datasets? – yes
  4. 4.
    If yes, please complete the list below.
    • A motivation is given for why the experiments are conducted on the selected datasets – yes
    • All novel datasets introduced in this paper are included in a data appendix – partial (cannot share raw video data containing identifiable patient information)
    • All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes – partial (cannot share raw video data containing identifiable patient information)
    • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations – NA
    • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available. – NA
    • All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisficing. – NA
  5. 5.
    Does this paper include computational experiments? – yes
  6. 6.
    If yes, please complete the list below.
    • Any code required for pre-processing data is included in the appendix – yes
    • All source code required for conducting and analyzing the experiments is included in a code appendix. – yes
    • All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. – yes
    • If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results. – yes
    • This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks. – yes
    • This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics. – no (evaluation metrics are widely used)
    • This paper states the number of algorithm runs used to compute each reported result. – yes
    • Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information. – yes
    • The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank). – yes
    • This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments. – yes (in the extended version)
    • This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting. – yes (in the extended version)